1 I am trying to solve the differential-algebraic equations (DAEs) which arise from the electrical transient stability problem, by using parallel computing. This work is associated with the NPAC project, “High Performance Computing and Parallel Processing in the Utility Industry.” I am in the process of writing a small electrical transient stability simulation program for testing and benchmarking sequential and parallel DAE solvers.

2 I expect to continue this work in phase II of the same NPAC project.

3 I am the coauthor of two conference papers.

Anthony Skjellum, Alvin Leung, Steven G. Smith, Robert D. Falgout, Chares H., Still and Chuck H. Baldwin, “The Multicomputer Toolbox – First-Generation Scalable Libraries,” HICSS-27, June, 1993. Alvin Leung, Anthony Skjellum and Geoffrey Fox, “Concurrent DASSL: A Second-Generation, DAE Solver Library,” Conference on Scalable Parallel Libraries, Oct., 1993.

4 My thesis will be related to algorithms for solving differential-algebraic equations on parallel computers. My academic advisor at Syracuse is Prof. Geoffrey Fox. My other thesis advisors are Sanjay Ranka and Anthony Skjellum.